### v1

# Exploring Context-aware and LLM-driven Locomotion for Immersive Virtual Reality

Abstract
Locomotion plays a crucial role in shaping the user experience within virtual reality environments. In particular, hands-free locomotion offers a valuable alternative by supporting accessibility and freeing users from reliance on handheld controllers. To this end, traditional speech-based methods often depend on rigid command sets, limiting the naturalness and flexibility of interaction. In this study, we propose a novel locomotion technique powered by large language models (LLMs), which allows users to navigate virtual environments using natural language with contextual awareness. We evaluate three locomotion methods: controller-based teleportation, voice-based steering, and our language model-driven approach. Our evaluation measures include eye-tracking data analysis, including explainable machine learning through SHAP analysis as well as standardized questionnaires for usability, presence, cybersickness, and cognitive load to examine user attention and engagement. Our findings indicate that the LLM-driven locomotion possesses comparable usability, presence, and cybersickness scores to established methods like teleportation, demonstrating its novel potential as a comfortable, natural language-based, hands-free alternative. In addition, it enhances user attention within the virtual environment, suggesting greater engagement. Complementary to these findings, SHAP analysis revealed that fixation, saccade, and pupil-related features vary across techniques, indicating distinct patterns of visual attention and cognitive processing. Overall, we state that our method can facilitate hands-free locomotion in virtual spaces, especially in supporting accessibility.
keywords:
virtual reality, large language models, locomotion, eye tracking.Introduction
Overview of our LLM-driven locomotion system in VR.
With the rapid advancement of virtual reality (VR) technologies, VR systems have started to be used widely in various domains and purposes, such as education [33], entertainment [5], healthcare [35] and training [79]. The quality of user interaction within immersive environments is essential to shape user experience positively. One of the important aspects of user experience is locomotion, which refers to how users move within virtual environments. Locomotion affects many factors including user immersion [12], task performance [26], overall comfort [39], and more importantly the cybersickness [25]. Among various techniques, teleportation by using handheld controllers remains one of the most widely adopted methods [62].
However, hands-free locomotion is particularly important for scenarios where users must keep their hands free, such as during multitasking or for applications focusing on accessibility. To address this, researchers proposed different locomotion techniques with varying input modalities, including voice [31, 20, 39], gaze [23], and gesture [49]. Among those, voice-based systems provide a natural and intuitive alternative, with techniques that include continuous voice-based steering [39] and teleportation methods controlled through verbal destination input [31, 20]. However, those previous approaches mainly depend on predefined command sets and rule-based mappings, which require users to speak unnaturally. This process restricts the flexibility and intuitiveness of the interaction experience, which can negatively affect user satisfaction, usability, and immersion.
Recent advancements in large language models (LLMs) such as GPT-4 offer new opportunities for developing more natural and intelligent voice-based interaction systems in VR [15]. These models can accurately interpret user instructions expressed in natural language in VR and effectively identify intent, particularly when the environmental context is incorporated into the prompt. Moreover, as they support multiple languages, it is possible to enable multilingual interaction within the same system setup and make virtual environments more generic with minimal effort while supporting user diversity.
In this study, we propose a locomotion method that utilizes an LLM to enable natural language-based, hands-free navigation in VR by addressing key limitations of existing voice-based techniques. Steering-based voice locomotion typically relies on continuous movement, which can cause discomfort and requires users to speak using rigid, predefined commands. While teleportation-based voice methods reduce motion sickness through instant movements, they still depend on rule-based grammar structures and often involve manual object labeling or predefined mappings for speech recognition and semantic understanding. In contrast, our approach allows users to express free-form spoken instructions without memorizing fixed commands or grammar structures. The system interprets these instructions in real time using contextual information dynamically extracted from the virtual environment, such as object names, colors, and positions, enabling flexible, intuitive, and scalable navigation without needing handcrafted rules or annotations.
To evaluate the effectiveness of our method, we conducted a user study in a town-like virtual environment, comparing three locomotion techniques: teleportation, voice-based steering through fixed commands, and our LLM-driven locomotion. We collected user feedback through standardized questionnaires and behavioral data using eye tracking. The questionnaires measure usability through SUS [16], presence through IPQ [71], cognitive load through NASA-TLX [36], and cybersickness through CSQ-VR [45]. The eye-tracking analysis provided insights into cognitive behavior, visual attention, and engagement. Furthermore, we trained machine learning models to classify the locomotion technique based on eye-tracking features and explored how each feature contributed to the models’ decisions by applying explainable artificial intelligence using SHAP analysis [50]. This process helped us identify the most influential gaze features in different locomotion techniques. Our findings indicate that teleportation is the fastest locomotion technique, as expected, while the LLM-driven approach enables slightly faster completion times than the voice-based steering method, with greater improvement observed in the second half of the task, indicating increased adaptability. In terms of user experience, the LLM-driven locomotion technique offers a similar level of usability to voice-based steering while improving spatial presence and maintaining low levels of cybersickness. Eye-tracking analysis revealed that the LLM-driven method resulted in fewer but longer fixations and shorter saccade durations, indicating more focused visual engagement compared to continuous movement. Additionally, SHAP analysis showed that gaze behavior varied across locomotion techniques, highlighting fixation duration, saccade dynamics, and pupil diameter as key features that differentiated users’ visual responses depending on how they navigated the virtual environment. Therefore, this work identifies LLM-driven locomotion as a hands-free and natural alternative to traditional techniques.
1 Related Work
This section provides previous works in three folds, including locomotion techniques, hands-free interaction, and LLMs in VR.
1.1 Locomotion in VR
Locomotion is one of the main components of user interaction in virtual environments, as it enables users to navigate and engage with digital spaces beyond their immediate physical boundaries. Without an effective locomotion mechanism, users are limited to static or highly constrained experiences, which may significantly diminish the immersive potential of virtual reality. Previous research proposed a broad spectrum of locomotion techniques in the literature [3, 53], and they are generally categorized into five types: walking-based, steering-based, selection-based, manipulation-based, and automated locomotion. Walking-based approaches simulate natural gait and include methods such as redirected walking [17, 46], omnidirectional treadmills [77], walking-in-place [75, 30], and arm swinging [78]. Steering-based techniques allow users to control their movement direction through inputs such as joysticks [68, 26], gaze or head orientation [63, 10], or voice commands [39], facilitating continuous movement through virtual spaces. Selection-based locomotion enables users to select a destination and instantly or gradually transport them to the destination. Among these, teleportation [11] is one of the most widely adopted techniques, typically involving instant transportation, while others support smooth transitions toward selected points [48, 56]. Furthermore, manipulation-based methods allow users to directly control their virtual position by interacting with the scene or camera system [12, 22, 65]. In contrast, automated locomotion involves moving users along predefined routes or sequences under system control, often with minimal user input [61, 66].
In the context of voice-based teleportation-like locomotion for hands-free navigation, Ferracani et al. [31] proposed a system for immersive museum experiences that enabled users, particularly those with motor impairments, to navigate by issuing semantic voice commands such as “I would like to see The Starry Night.” A rule-based grammar and ontological reasoning system interpreted these commands; however, the system required a manual definition of both object metadata and grammar rules, limiting its flexibility and scalability in more dynamic environments. Similarly, Calandra et al. [21] evaluated speech-based navigation methods for training scenarios, including voice-only, voice with gaze, and hybrid approaches. While effective, their techniques relied on predefined destination names or required disambiguation logic through user interface panels and structured grammar templates, which may reduce naturalness and increase user cognitive load.
1.2 Hands-Free Interaction in VR
Hands-free interaction is an essential feature of immersive VR systems [57], particularly in scenarios where users must keep their hands free for primary tasks such as surgery or equipment operation [32]. Monteiro et al.[58] conducted a comprehensive systematic review of hands-free interaction techniques in immersive VR, categorizing modalities based on input sources such as voice, eye gaze, and head movement. Among these, voice remains the most extensively studied modality due to its integration into commercial head-mounted displays (HMDs) and advancements in voice processing technologies [29, 31].
Different voice-based interaction techniques exist, including command-based and natural language processing (NLP)-based techniques. Command-based techniques operate on a limited set of predefined phrases (e.g., “go forward,” “stop”), which improves the command recognition accuracy and allows for offline or lightweight computational processing [70]. These techniques are efficient and easy to learn, as users only need to memorize a small set of commands. However, their rigidity can be limiting, and they are not well-suited for complex tasks or dynamic interfaces, as they can be perceived as unintuitive by users. In contrast, NLP-based techniques enable users to issue free-form, natural language commands (e.g., “rotate by 45 degrees” or “select the purple circle”), offering more intuitive and expressive interaction [37]. These techniques are well-suited for symbolic input, parameterized commands, and narrative-driven tasks. However, they often depend on rule-based mechanisms that are sensitive to recognition errors caused by accent variation, phrasing differences, or background noise. Other voice-based approaches include symbolic input through voice-to-text [2] and non-verbal vocal cues [74], expanding the range of interaction possibilities.
Beyond voice, researchers have also utilized other hands-free modalities, including eye gaze for target selection or interaction with user interfaces [52, 9], often in combination with other modalities to mitigate issues such as the “Midas Touch” effect [42]. Head gaze, facilitated by HMD tracking, is another intuitive input method that supports pointing and navigation [21]. More novel modalities, such as facial expressions [24], body postures [34], and brain-computer interfaces (BCIs) [51], have gained attention for their potential to enhance accessibility and personalization. However, these approaches often require special hardware or user training, which can limit their widespread adoption.
1.3 LLMs for VR
The versatile computational capabilities of foundation models and LLMs make them suitable for handling complex tasks across different domains, such as medicine [4] and education [41]. Considering the fact that LLMs can personalize user experiences often with few prompts and do not require extensive manual labor, such as in the form of annotations, similar to many domains, the VR community has also started to integrate those models into their workflows. In fact, integration of LLMs can facilitate inclusion and equity in virtual spaces and help increase users’ engagement as they can provide infrastructure for intuitive interaction experience [15]. To this end, Buldu et al. [18] focused on delivering a speech-based interaction framework that relies on LLMs in VR by combining different speech-to-text, text-to-speech, and language models. Furthermore, Lau et al. [47] argued for using LLMs for personalization in VR spaces and found that personalization through LLMs in VR boosts engagement and learning interest, indicating the potential benefits of LLMs to support users. Furthermore, De La Torre et al. [27] utilized LLMs to create interactive and virtual spaces, especially for producing and editing objects and scenes. The authors showed the effectiveness of their framework through a usability study in which participants provided positive feedback. This wide range of use cases of LLMs in VR shows their potential to make VR more engaging and accessible for users. In the context of this work, when hands-free locomotion is considered, one of the most straightforward ways to obtain the user’s input is through voice input [39]. However, previous research often considers fixed voice-based commands, limiting the user inputs’ generalizability. We address this issue with LLMs by incorporating contextual information from the scene to facilitate hands-free and context-aware locomotion in VR.
2 Methodology
In this study, we investigated the effectiveness of three distinct locomotion techniques in VR: teleportation, voice-based steering through fixed commands [39], and our proposed LLM-driven locomotion approach. To evaluate these techniques, we designed a virtual environment where participants were given a task that required them to navigate through the scene in our between-subjects design user study. We collected both objective data, including eye tracking to assess attention and cognition, and subjective data, including the System Usability Scale (SUS), the Igroup Presence Questionnaire (IPQ), the Cybersickness in Virtual Reality Questionnaire (CSQ-VR), and the NASA Task Load Index (NASA-TLX). These questionnaires measured usability, sense of presence, cybersickness, and cognitive load. In the following subsections, we describe the locomotion techniques and explain the details of our user study. The Institutional Review Board of [Blind Institute] approved our study protocols and data collection procedures.
2.1 Locomotion Techniques
In this study, we implemented three different locomotion techniques, including teleportation, voice-based steering through fixed commands, and proposed LLM-driven locomotion. We treat the teleportation and voice-based steering as general and hands-free baselines, respectively.
2.1.1 Controller-based Teleportation
We use controller-based teleportation as the baseline locomotion method, representing a conventional and widely adopted technique in VR [11, 3]. This method is typically perceived as easy to learn and use, and we chose it because of its easy adaptability to most virtual environments. Users are generally comfortable with this locomotion technique [62]. In this method, the user points to a location using the controller button and is instantly transported to the selected position. In the default implementation, a curved line is displayed to guide the user. A green arc indicates a valid teleportation target, while a red arc signals that teleportation to the selected area is not possible. When a valid location is selected, the user is immediately transported to this location after participants pull their fingers from the button. This method minimizes motion-induced discomfort and is generally well-tolerated by users [11].
2.1.2 Voice-based Steering through Fixed Commands
We utilize a voice-based steering locomotion method based on fixed commands, which is one of the most preferred hands-free techniques by users [39]. It enables continuous directional control using a predefined set of voice commands. We selected this technique primarily because of its generic commands, which provide high adaptability across diverse virtual environments without requiring significant customization. In contrast, other voice-based locomotion methods often rely on environment-specific configurations or require extensive manual annotations, which limits their scalability in large or dynamic virtual environments. For instance, traditional landmark-based approaches require semantic mapping and labeling for all recognizable objects in the scene, demanding significant manual effort. Similarly, number-grid-based systems [39] become increasingly impractical as the environment grows in size, as they would require an overwhelming number of grid identifiers to cover the entire space.
In this technique, we utilized the VOSK speech recognition engine [73] due to its robustness in handling a variety of voice commands and its demonstrated reliability in previous applications. Participants controlled movement through a predefined set of voice commands, including <go forward>, <go back>, <turn left>, <turn right>, and <stop>. Additional commands such as <faster> and <slower> are also supported, allowing users to adjust their walking speed dynamically across four discrete levels: 1.4 m/s, 2.8 m/s, 4.2 m/s, and 5.6 m/s. The baseline value of 1.4 m/s aligns with the average human walking speed [59], ensuring comfort, while higher speeds enable more efficient exploration for experienced users [23, 19].
Although it is mainly designed as a hands-free alternative for our experiment setting to minimize any errors or effects of design preferences, we use a trigger button to start talking and use the same button to stop recording to minimize the failures in voice recognition. Once a valid command such as <go forward> is provided, the user starts moving continuously at a constant speed. While directional changes could be controlled using the <turn left> and <turn right> commands, users are also free to rotate naturally using their physical head and body movements. Nevertheless, unlike teleportation, which involves instantaneous movement and minimizes sensory conflict, voice-based steering relies on continuous movement, which has been shown to increase the likelihood of VR-induced motion sickness [39]. While teleportation kind instant movement-based methods do not trigger expectations of physical motion, continuous movement generally introduces a greater mismatch between visual and vestibular signals, leading to increased discomfort [55].
2.1.3 LLM-driven Locomotion
Our proposed LLM-driven locomotion method introduces a novel approach to hands-free, natural language-based teleportation, enabling intuitive navigation in virtual environments through free-form voice commands. In contrast to conventional voice-based locomotion techniques [39, 21, 31], which rely on rigid command sets or rule-based grammar structures, our approach leverages LLMs to interpret open-ended and flexible instructions through contextual understanding. This approach eliminates the need for users to memorize specific commands or syntax, resulting in a more natural and engaging interaction experience. Moreover, the method offers inherent multilingual support without requiring separate command definitions for each language, thereby improving both accessibility for users and scalability for developers. To navigate the environment, users can employ naturally phrased voice commands, such as <go to the red house> or <move 50 meters forward>, which the system processes contextually to determine the intended destination. Thus, our LLM-driven approach can be considered an advanced, dynamically context-aware extension of landmark-based teleportation [39, 31, 21], as it eliminates the need for predefined object mappings by leveraging real-time scene understanding.
To support this, we utilized the pipeline illustrated in Figure Exploring Context-aware and LLM-driven Locomotion for Immersive Virtual Reality, which takes the system prompt, the user’s voice input, and contextual scene information to construct a comprehensive user prompt. The system prompt primarily describes the purpose of the agent as a “navigation assistant in Unity” and explains the structure of the user prompts, specifying the rules and expected output format. It also provides three examples. The user prompt includes the user’s command and contains explicit environmental context, including the positions of nearby landmarks, buildings, and vehicles currently visible to the user. The LLM processes this information to generate precise target coordinates. Subsequently, these coordinates are dynamically mapped to the closest valid navigable location within the virtual environment based on predefined walkable areas. In scenarios where the generated coordinates correspond to multiple possible valid locations, such as intersections or corners, the system employs the user’s gaze direction to select the most appropriate position by calculating the smallest angle between the gaze vector and available street directions to align with the user’s intended movement trajectory. Our system primarily relies on eye gaze for this calculation; however, head gaze is used as a fallback if eye gaze data is unavailable or unreliable. Additionally, to ensure robustness, the system incorporates several fallback mechanisms. If the LLM returns excessive text along with the location rather than only the expected coordinate, we dynamically extract the coordinate from the response. If the LLM fails to generate meaningful or interpretable output, or if the user’s command is ambiguous, the user’s position remains unchanged, effectively preventing unintended movement. Successful commands are visually indicated to the user by a curved line marking the calculated destination point, and after a two-second interval, the user is transported to the new location. Similar to the voice-based steering through fixed commands, we use a trigger button to start and stop voice recording during the experiment to minimize the effect of design preferences.
2.2 User Study
2.2.1 Participants
The study included data from 63 participants; however, due to technical issues, three participants were excluded, resulting in a final sample of 60 participants (20 per locomotion condition). The mean age is 27.1 years (SD = 5.33). The gender distribution is 60% women (), 36.67% men (), 1.67% non-binary (), and 1.67% who preferred not to disclose (). In terms of educational background, 46.67% hold a bachelor’s degree or equivalent (), 33.33% had a master’s degree or equivalent (), 13.33% had completed high school or equivalent (), and 6.67% hold a doctorate or equivalent (). Most participants were students (73.33%, ), while others were employed (20.00%, ), unemployed (5.00%, ), or categorized as other (1.67%, ). 98.33% () of participants had used LLMs before. Regarding VR experience, 70.00% () had used VR previously, although only 8.33% () own a VR device.
2.2.2 Apparatus
We designed the environment in Unity and implemented three different locomotion techniques: teleportation using the VIVE controller, fixed command-based locomotion [39] using the VOSK speech recognition engine [73], and LLM-driven locomotion utilizing the open-source CUIfy package [18], which provides an optimized pipeline for LLM-based interaction through speech-to-text models. For speech recognition, we used a locally hosted medium Whisper model [64], as it offers greater robustness compared to smaller versions, even though this introduced higher latency. For language understanding, we employed ChatGPT-4o [60] as the core LLM. We ran the experiments on a Varjo XR-3 [76] HMD, connected to a desktop system equipped with an Intel Core i7-13700K processor, 32 GB of RAM, and an NVIDIA GeForce RTX 4080 GPU. We collected eye-tracking data with the XR-3’s built-in eye tracker at a 200 Hz sampling rate.
2.2.3 Experimental Design
We employed a between-subjects design, where each participant experienced only one of the three locomotion methods: teleportation, fixed voice command-based locomotion, or LLM-driven locomotion. This approach ensured that individual learning effects, fatigue, or cybersickness did not influence performance across different conditions. All participants navigated the same virtual environment. We conducted the study in a VR environment developed using Unity, designed to resemble a virtual small town with a simple street layout consisting of four vertical and four horizontal roads, as shown in Figure 1. We designed a two-step navigation task to evaluate each locomotion method within a task-oriented context. We first instructed participants to locate a bank in the virtual town. Each participant started at the same position, marked by the blue “X” in Figure 1, and completed the task by finding the target locations at identical coordinates, indicated by pink holograms on the street. To make the task intuitive, we placed a prominent purple hologram in front of the bank, making it easy to identify. Upon reaching the bank, the first hologram disappeared, and a second one appeared at the town’s exit as a second target. Participants were then required to find and reach this exit point to complete the navigation task.
2.2.4 Procedure
We first welcomed the participants and asked them to provide written consent to participate in the experiment. Next, we explained the (randomly) assigned locomotion condition to them in detail. Participants then wore the HMD, and the session began with a demo scene designed to help them become familiar with the locomotion technique and VR. Eye-tracking calibration was also performed during this phase. Once participants felt comfortable about the VR and controls, following the demo, the main experiment commenced. As instructed, participants were supposed to navigate the virtual environment and reach two holograms placed at specific locations. After completing the task, participants completed questionnaires, including the NASA-TLX, SUS, IPQ, CSQ-VR, and a demographic survey. Later, they were compensated with €7.5 for an experimental session that was expected to last half an hour in total.
2.2.5 Measurements
We conducted a comprehensive analysis of each locomotion technique using statistical methods. We collected eye-tracking data to objectively evaluate participants’ engagement and cognitive behavior. In addition, we collected their feedback through standardized questionnaires, including the SUS, IPQ, CSQ-VR, and NASA-TLX. We describe these details in the following.
Eye Tracking Analysis
To evaluate the locomotion techniques, we conducted an eye-tracking analysis, as eye tracking has been widely used to assess user engagement, attention, behavior, and cognitive processes during interaction [38, 80]. We classified samples as either fixations or saccades from the raw eye-tracking data. Fixations represent moments when the eyes remain steadily focused on a specific point, while saccades are rapid eye movements between fixation points. We analyzed several features to understand users’ visual attention and cognitive states. The fixation rate, which reflects the frequency of attentional shifts, is commonly used to assess how users allocate their attention during a task [28]. Mean fixation duration, which represents the length of time users dwell on specific points, has been linked to deeper cognitive processing and enhanced information encoding [43]. The mean saccade duration captures the speed of gaze transitions, indicating the efficiency of visual search behavior. Pupil diameter provides insights into the mental effort and arousal, correlating with the cognitive workload, fatigue, and engagement [6, 40]. Additionally, we identified blinks and extracted both blink frequency and duration, as these factors are associated with cognitive load and visual fatigue [7]
System Usability Scale (SUS)
To assess usability, we employed the System Usability Scale (SUS) [16], one of the most widely used tools for evaluating system usability. The questionnaire consists of 10 items, with participants rating each statement on a 5-point Likert scale, ranging from 1 (strongly disagree) to 5 (strongly agree). The SUS score is calculated by summing the scores for each question and then scaling the total, resulting in a final score ranging from 0 to 100, where higher scores indicate better usability.
Igroup Presence Questionnaire (IPQ)
To assess the sense of presence in the virtual environment and the influence of each locomotion technique on it, we used the Igroup Presence Questionnaire (IPQ) [72]. While based on the theoretical framework introduced by Schubert et al. [71], the IPQ consists of 14 questions designed to evaluate presence. It measures four dimensions: Spatial Presence (SP), Involvement (INV), Experienced Realism (REAL), and General Presence (PRES). Each question is rated on a 7-point Likert scale, and the mean score is calculated for each category.
CyberSickness in Virtual Reality Questionnaire (CSQ-VR)
To evaluate cybersickness, we used the Cybersickness in Virtual Reality Questionnaire (CSQ-VR) [45], which is designed to assess both the intensity and type of cybersickness symptoms, including nausea, disorientation, and oculomotor discomfort, experienced during VR exposure. The questionnaire comprises six items, each rated on a 7-point Likert scale.
NASA-Task Load Index (NASA-TLX)
To assess the impact of the locomotion technique on participants’ cognitive load, we employed the NASA-TLX [36], a widely used tool for evaluating perceived workload across multiple dimensions.
| Event | Velocity () | Duration () |
| Fixation | ||
| Saccade |
2.2.6 Data Processing
To analyze eye-tracking data and pupillometry data, we first need a multi-step data processing pipeline. First, based on gaze velocity, we classified eye movement data into fixations and saccades using the Velocity-Threshold Identification (I-VT) algorithm [67, 44]. Stable gaze movements are classified as fixations, while faster movements are labeled as saccades. We also incorporated head movement data to improve the fixation detection reliability. We only considered a fixation when both the gaze and head remained stable, similar to previous works [33, 1]. We present the detailed criteria for classifying fixations and saccades—including velocity and duration thresholds in Table 1.
To ensure an accurate interpretation of pupil diameter, we applied a Savitzky-Golay filter [69] to smooth the raw signal and reduce short-term noise. Subsequently, we performed a divisive baseline correction using a one-second pre-stimulus interval [54], allowing for normalization across participants and experimental trials. Pupil diameter was used as an indicator of objective cognitive load and engagement, similar to prior virtual reality studies [33, 14, 13].
Additionally, we analyzed blinks using the Varjo XR-3’s eye openness data. We identified a blink when the eye openness value reached zero, following a consistent decrease in eye openness, indicating a natural eyelid closure. To minimize false positives caused by tracking loss, we discarded blink candidates that did not exhibit a decreasing eye openness trend before reaching zero.
2.2.7 Analysis
We conducted a series of statistical analyses on both questionnaire responses (e.g., SUS, NASA-TLX) and eye-tracking measures to examine differences across experimental conditions. For each dependent variable, we performed a one-way Analysis of Variance (ANOVA) to evaluate the effect of the locomotion technique (teleportation, LLM-driven locomotion, and voice-based steering). When the ANOVA indicated a statistically significant effect, we used Tukey’s Honest Significant Difference (HSD) post-hoc test to determine which pairs of conditions differed significantly. This approach controlled the family-wise error rate while comparing all condition pairs.
We checked the assumptions of normality and homogeneity of variance before conducting ANOVA, applying Shapiro-Wilk tests for normality and Levene’s test for homogeneity of variances. In cases where these assumptions were violated, we conducted the non-parametric Kruskal-Wallis H test as an alternative to ANOVA. We conducted all statistical tests using . Where relevant, we also report the effect sizes (e.g., ) to support the interpretation of findings.
2.2.8 Model Building
We conducted detailed analyses of eye-tracking data to investigate how different locomotion techniques influence user behavior. These analyses aim to reveal potential cognitive and attentional differences among participants using teleportation, voice-based steering locomotion, and LLM-driven locomotion techniques. Additionally, we trained classification models to predict the locomotion condition based on eye-tracking features. This process enables us to identify which gaze-related features are most affected by the locomotion technique and which features are the most predictive or distinctive in differentiating between the locomotion techniques. We provide the full list of features used in Table 2.
| Features | Statistical Metrics |
| Number of fixations | Total number of fixations |
| Fixation duration | Mean, Std, Min, Max, Sum of the fixation durations |
| Number of saccades | Total number of saccades |
| Saccade duration | Mean, Std, Min, Max, Sum of the saccade durations |
| Saccade peak velocity | Mean, Std, Min, Max of the saccade peak velocities |
| Saccade amplitude | Mean, Std, Min, Max of the saccade amplitudes |
| Sac-fix duration ratio | Ratio between saccade and fixation total durations |
| Sac-fix count ratio | Ratio between saccade and fixation counts |
| Number of blinks | Total number of blinks |
| Blink duration | Mean, Std, Min, Max of the blink durations |
| Pupil diameter | Mean, Std, Min, Max of the normalized pupil diameters |
We evaluated the performance of five classifiers: Random Forest, LightGBM, k-Nearest Neighbor (k-NN), Support Vector Machine (SVM), and Gradient Boosting. The classification was performed on data from 20-second windows, with 80% of the data used for training and 20% for testing. Hyperparameters were optimized using a stratified 5-fold cross-validation strategy to ensure robust model performance across classes. To understand the models’ decision-making processes, we applied SHAP (SHapley Additive exPlanations) [50] to interpret feature importance in a transparent and model-agnostic fashion. SHAP values provide a detailed attribution of feature contributions toward classification outcomes, helping us identify which specific aspects of gaze behavior, such as fixation duration, saccade amplitude, or pupil diameter, are most influenced by the underlying locomotion technique.
3 Results
3.1 Performance Analysis
Figure 2 presents the participants’ task completion times across the three locomotion conditions. On average, participants completed the task in , using teleportation, , with the LLM-driven locomotion, and , with the fixed voice command approach. As expected, teleportation was the fastest method, serving as the baseline for comparison. A one-way ANOVA revealed a significant effect of locomotion condition on task completion time, , , with a large effect size (). Post-hoc tests showed teleportation was significantly faster (), while the LLM-driven and fixed command methods showed no significant difference (), indicating comparable performance between the hands-free methods.
We further analyzed the time required to reach each of the two sequential targets to explore potential learning effects. For the first target (the bank in the virtual town), participants in the voice-based steering condition reached it in , , while those in the LLM-driven condition took , . From the bank to the second target, the LLM-driven condition completed it in , , whereas the voice-based steering condition required , . Although a tendency toward improvement was observed in both conditions, no statistically significant differences were found.
In addition to completion times, we also measured system latency for LLM-driven locomotion, as teleportation and voice-based steering were nearly instantaneous. The speech-to-text component, using a locally hosted medium-sized Whisper model, had an average processing time of , , while the LLM component, utilizing the ChatGPT-4o API, had an average response time of , . In total, the average end-to-end processing time remained under 1.5 seconds (, ).
3.2 Eye-Tracking Analysis
Fixation rate
To assess cognitive engagement under different locomotion techniques, we analyzed the fixation rate, defined as the number of fixations per second, as shown in Figure 6. ANOVA revealed a statistically significant effect of locomotion type on fixation rate, , , with a large effect size (), indicating that the type of locomotion significantly influenced visual attention patterns. Additionally, the statistics indicated that the voice-based steering condition resulted in the highest fixation rate (, ), followed by teleportation (, ), and voice LLM (, ). Post-hoc comparisons using Tukey’s Honest Significant Difference (HSD) test showed that both teleportation and voice LLM conditions had significantly lower fixation rates compared to the voice-based steering condition ( for both comparisons). However, no statistically significant difference was observed between the teleportation and voice LLM conditions (), suggesting similar visual engagement across these two hands-free techniques.
Mean Fixation Duration
Mean fixation durations are presented in Figure 6, and all values are reported in milliseconds. ANOVA revealed no statistically significant differences between conditions, , . Among the conditions, teleportation showed the highest mean fixation duration (, ms), followed by LLM-driven locomotion (, ms), and voice-based steering with the lowest mean duration (, ms).
Mean Saccade Duration
Mean saccade durations for each condition are presented in Figure 6, and all values are reported in milliseconds. ANOVA revealed a statistically significant effect of locomotion type on mean saccade duration, , , with a large effect size (). Descriptive statistics showed that the voice-based steering condition had the highest mean saccade duration (, ms), followed by teleportation (, ms), and the LLM-driven condition with the lowest duration (, ms). Post-hoc tests showed mean saccade duration was significantly lower in the LLM-driven condition than in teleportation () and voice-based steering (), with no difference between the latter two ().
Pupil Diameter
We analyzed the mean baseline-corrected pupil diameter across the three locomotion conditions to assess participants’ cognitive load as shown in Figure 6. ANOVA revealed no statistically significant differences between the conditions, , . Although not statistically significant, the descriptive statistics suggest a trend. The teleportation condition caused the lowest mean pupil diameter (), while the voice-based conditions showed higher and relatively similar values (Voice LLM: ; Voice-Based Steering: ). This pattern indicates slightly higher cognitive effort in the voice-based hands-free methods compared to teleportation.
3.3 Questionnaire Analysis
3.3.1 CSQ-VR Questionnaire
Figure 7 illustrates the CSQ-VR scores across locomotion techniques, covering overall cybersickness as well as the subcategories of nausea, vestibular, and oculomotor symptoms. The general score range for the CSQ-VR questionnaire is between 0 and 12 for nausea, vestibular, and oculomotor symptoms, and between 0 and 36 for overall sickness. Overall, participants reported low levels of cybersickness across all conditions, with no statistically significant differences observed. Teleportation showed the lowest scores in overall sickness (, ), nausea (, ), and vestibular discomfort (, ), suggesting a generally more comfortable experience. For overall sickness, the LLM-driven condition showed the highest scores (, ); however, the voice-based steering condition resulted in the highest levels of nausea (, ) and vestibular symptoms (, ), while the LLM-driven condition fell between the two. Interestingly, in the oculomotor discomfort, the voice-based steering condition showed the lowest scores (, ), whereas both teleportation (, ) and the LLM-driven condition (, ) reported higher values. No statistically significant differences were found between conditions (), and we report the overall trends.
3.3.2 IPQ Scores
The IPQ results, shown in Figure 8, present scores across four dimensions of presence: general presence, spatial presence, realism, and involvement, for each locomotion condition. Scores range from 0 to 6. The LLM-driven locomotion method received the highest ratings in both general presence (, ) and spatial presence (, ), indicating that participants experienced a stronger sense of immersion in the virtual environment. In the realism dimension, the voice-based steering condition scored the highest (, ), suggesting that participants perceived the environment as more realistic when using voice command-based steering. For involvement, teleportation resulted in the highest score (, ), while the LLM-driven locomotion (, ) and voice-based steering methods (, ) showed similar values. We did not find statistically significant differences across the conditions ().
3.3.3 NASA-TLX Questionnaire
The NASA-TLX results, illustrated in Figure 10, present the perceived cognitive load associated with each locomotion technique. The score range for NASA-TLX is from 0 to 100. In general, participants did not report experiencing substantial cognitive load during the experiments. The teleportation condition yielded the lowest average score (, ), indicating the least cognitive demand among the three methods. In contrast, the LLM-driven locomotion method showed the highest average cognitive load (, ), while the voice-based steering method fell in between (, ). Despite these slight differences in mean values, statistical analysis revealed no significant differences between the distributions across conditions().
| Model | Accuracy |
| Random Forest | 0.6769 |
| LightGBM | 0.7154 |
| k-NN () | 0.6462 |
| SVM (RBF kernel) | 0.5692 |
| Gradient Boosting | 0.6846 |
3.3.4 SUS Questionnaire
The SUS results, shown in Figure 10, reflect perceived usability. While there was no statistically significant difference between the techniques (), teleportation, the baseline method, received the highest usability score (, ), indicating a high level of user satisfaction and ease of use. Both the LLM-driven (, ) and voice-based steering conditions (, ) scored slightly lower. These results suggest generally good usability for both as hands-free alternatives, though less favorable than teleportation.
3.4 Model Evaluation and Feature Importances
We conducted a comprehensive analysis to evaluate the performance of various machine learning models and identify key eye-tracking features changing with the locomotion method. The baseline accuracy obtained by predicting the majority class fix command-based voice steering was . In the test set, the class distributions are 55 samples () for LLM-driven locomotion, 56 samples () for fix-command-based voice steering, and 19 samples () for teleportation. We summarize the classification accuracies of each model in Table 3. Among these, LightGBM achieved the highest performance, with an accuracy of , followed by Gradient Boosting and Random Forest.
To interpret the effect of locomotion techniques on eye movement characteristics, we used SHAP to evaluate the contribution of individual features within the LightGBM classifier. The SHAP summary plot is shown in Figure 11, illustrating the average impact of each feature on the model’s output, aggregated by classes teleportation, voice-based steering through fixed commands, and LLM-driven locomotion. According to the SHAP analysis, total fixation duration, minimum pupil size, and mean saccade duration emerged as the most important features for distinguishing between locomotion techniques. These features are commonly linked to cognitive load, as well as visual attention, engagement, and visual search behavior, suggesting that user interaction patterns varied across conditions. Additionally, total saccade duration and maximum blink duration contributed substantially to the model’s predictions, indicating further differences in attentional strategies. Other influential features included the standard deviation of saccade duration, minimum saccade peak velocity, and blink count. Overall, the results indicate that fixation behavior, saccade dynamics, blink activity, and pupil-based measures collectively shape the gaze responses associated with each locomotion technique, reflecting distinct patterns of cognitive processing and user engagement in VR.
4 Discussion
We investigated three locomotion techniques within a VR environment: teleportation, fixed voice command-based locomotion and context-aware LLM-driven locomotion. While teleportation is used as a baseline condition, the primary focus was to evaluate the effectiveness of the two hands-free techniques in terms of usability, user comfort, cognitive demand, and immersive experience.
4.1 Performance and User Experience
Teleportation, while the fastest method, relies on hand controllers, making it unsuitable for hands-free scenarios. Among the hands-free methods, voice-based steering initially enabled faster navigation to the first target, but participants quickly adapted to the LLM-driven method, resulting in more efficient performance in the second phase and faster overall completion. This learning effect suggests that users quickly adapted to the natural language system and were able to make effective use of its flexibility, despite being completely unfamiliar with this form of locomotion.
All three methods showed comparable results across usability, cognitive load, cybersickness, and presence, with no significant differences observed. These results indicate that both voice-based locomotion alternatives are effective for users and can be as helpful as teleportation in accessibility-focused situations in VR. Teleportation received the highest usability score indicating “excellent” usability. Both voice-based methods received scores between 75 and 80, reflecting “good” usability. These results demonstrate that both hands-free techniques were generally well-received by users, even though they did not match the efficiency or ease of teleportation.
The LLM-driven method, which enables instant transport to target locations, reduced nausea and vestibular discomfort compared to continuous movement in voice-based steering. This aligns with prior work showing that continuous motion is more likely to induce cybersickness [8]. Although all participants reached the same targets, those using teleportation completed the task much faster, resulting in significantly less exposure to the virtual environment—nearly four times shorter than other conditions. This reduced duration likely contributed to the lower overall cybersickness scores in the teleportation condition. Interestingly, the pattern reversed for oculomotor discomfort, where voice-based steering yielded slightly lower scores. This may indicate that its steady, continuous movement places less strain on the visual system than the abrupt position changes used in the other techniques.
Teleportation required the least cognitive effort, while the LLM-driven method showed slightly higher cognitive demand than voice-based steering. A similar trend was observed in the pupil diameter measurements. The increase in the LLM-driven condition may reflect the mental effort required to plan contextually relevant and efficient commands, which is not expected in teleportation or in command-based methods that rely on a simple and repeatable set of inputs. However, it may also promote deeper cognitive engagement and encourage more active exploration of the immersive virtual environment, potentially contributing to the heightened sense of presence reported in the LLM-driven locomotion condition. Additionally, although free-form commands may initially impose a higher cognitive load, this demand is likely to decrease with experience, which can be further investigated through longitudinal studies. Despite the slightly higher cognitive load associated with contextual commands, participants using LLM-driven locomotion reported the highest levels of general and spatial presence. Actively engaging with the environment to provide context-aware instructions likely strengthened their sense of being “present” compared to voice-based steering. In contrast, voice-based steering received the highest realism scores, possibly because its continuous walking pattern more closely mimics real-world locomotion.
In addition to the standardized questionnaires, participants provided positive qualitative feedback across all locomotion strategies, although each evaluated only one method. Teleportation was commonly described as intuitive and easy to adapt to, contributing to a smooth experience. Voice-based steering was perceived as natural due to its similarity to real-world walking and the feeling of continuous movement throughout navigation. However, participants occasionally attempted to use command variations beyond the predefined set, which sometimes led to unrecognized inputs. The LLM-driven approach received particularly positive feedback. Many participants described it as highly intuitive, engaging, and enjoyable to use. In fact, during the demo scene, some users experimented with different phrasings, suggesting that the open-ended input style encouraged curiosity, exploration, and deeper engagement with the virtual environment. However, there were a few instances where the system failed to understand commands, mainly due to limitations in the speech-to-text module.
4.2 Cognitive and Visual Attention
Voice-based steering resulted in significantly higher fixation rates, indicating increased visual scanning and attentional demands during continuous navigation. This result may be attributed to the nature of constant movement, where users need to actively monitor their direction and rapidly search the environment, often resulting in shorter and more frequent fixations. Similarly, mean fixation duration serves as an indicator of stable visual attention and cognitive engagement. Teleportation resulted in the longest fixations, followed by the LLM-driven method, while the voice-based steering condition showed the shortest durations. These eye movement patterns, particularly less frequent but longer fixations in the teleportation and LLM-driven conditions, suggest that users were able to plan their actions more deliberately and were generally aware of their intended destinations.
Mean saccade durations were significantly shorter for LLM-driven locomotion, suggesting more efficient and goal-directed gaze behavior, indicating that participants in the LLM-driven condition engage in more purposeful visual exploration. Additionally, while pupil diameter is primarily considered an indicator of cognitive load, increased pupil size may also reflect heightened engagement or arousal within the virtual setting. This finding is further supported by the higher spatial and general presence scores in the LLM-driven condition, suggesting that natural language interaction facilitated deeper cognitive engagement and immersion.
Classification models demonstrate the feasibility of using eye-tracking features to infer user interaction mode in real time, as these features varied across locomotion techniques. SHAP analysis indicated that fixation duration, saccade dynamics, and pupil diameter were key indicators for distinguishing user behavior. These distinct eye movement patterns, which reflect variations in visual attention and cognitive engagement, can serve as objective metrics for adaptive VR systems.
4.3 Accessibility and Practical Considerations
The LLM-driven locomotion method provides a highly accessible alternative for users with disabilities or temporary impairments, matching teleportation and fixed-command methods in usability and comfort. With its inherent multilingual support, users can interact naturally in their own languages without relying on language-specific phrasing. This eliminates the need for environment-specific adaptations or manual configuration, making the system highly adaptable and inclusive. Such flexibility is especially valuable in settings like medical training, education, and remote collaboration, where accessibility and ease of use are critical as VR adoption grows.
However, one challenge in implementing LLM-driven locomotion in real-world scenarios is the inherently dynamic and potentially unpredictable nature of LLM outputs, despite their strong performance across a wide range of conditions. Since they are not entirely controllable, ensuring consistent and predictable behavior may require additional error-handling mechanisms. These may include constraining the range of movement, validating target coordinates, or keeping the user in place when the model output is ambiguous or refers to non-navigable areas. Another challenge is finding the right balance between accuracy and performance in LLM and speech-to-text models. While reducing latency by using smaller models can improve responsiveness, it is crucial to maintain strong overall performance.
5 Conclusion
In this work, we proposed an LLM-driven, hands-free locomotion technique for VR that uses natural language instructions to enable instant teleportation, enhancing accessibility and addressing the motion sickness often associated with continuous voice-based steering. To assess the effectiveness of this method, we conducted a comparative evaluation of three locomotion techniques: teleportation, voice-based steering through fixed commands, and our LLM-driven approach. We used eye-tracking analysis to gain insight into users’ attention, engagement, and cognitive processes and complemented this analysis with standardized questionnaires, including the SUS, IPQ, CSQ-VR, and NASA-TLX, to assess usability, presence, cybersickness, and cognitive workload, respectively. While teleportation remains the most efficient locomotion method, it is often not suitable for hands-free scenarios. The proposed LLM-driven approach offers a promising alternative by combining natural language interaction with contextual awareness and instant movements while also inherently supporting multiple languages as an advantage. Although it requires slightly more cognitive effort, it enhances immersion and comfort, particularly in terms of reducing cybersickness. Overall, our results demonstrate that LLM-driven locomotion is a promising alternative for intuitive, accessible, and immersive hands-free navigation in VR environments. Future research could investigate personalization techniques, such as adapting LLM behavior based on user preferences or prior interactions, to enhance predictability and usability over time.